OpenStudy (sleepyjess):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (sleepyjess):

@sleepyhead314

OpenStudy (sleepyhead314):

(x-h)^2 = 4p(y-k) ?

OpenStudy (alexandervonhumboldt2):

The vertex is always halfway between the focus and the directrix, and the parabola always curves away from the directrix, so I'll do a quick graph showing the focus, the directrix, and a rough idea of where the parabola will go:

OpenStudy (sleepyjess):

yes, that is what they give me to use

OpenStudy (sleepyhead314):

*cheers* go alex go! :D cuz idk how the heck to do this xD

OpenStudy (alexandervonhumboldt2):

(y – (–2))^2 = 4(–1)(x – 1), or (y + 2)^2 = –4(x – 1)

OpenStudy (sleepyhead314):

you is copy pasting a lot @AlexandervonHumboldt2 ?

OpenStudy (alexandervonhumboldt2):

ok l;et me explain another way

OpenStudy (alexandervonhumboldt2):

Any point, (x0 , y0) on the parabola satisfies the definition of parabola, so there are two distances to calculate: Distance between the point on the parabola to the focus Distance between the point on the parabola to the directrix

OpenStudy (sleepyhead314):

@AlexandervonHumboldt2 would the vertex for this be (0, 0) ?

OpenStudy (alexandervonhumboldt2):

Distance between the point (x0 , y0) and (a , b): where (a, b) is the focus. Use distance formula here: sqrt((x0-a)^2+(y0-b)^2)

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

Distance between point ( x0 , y0) and the line y = c : =|y0-c|

OpenStudy (alexandervonhumboldt2):

Evaluate: sqrt((x0-a)^2+(y0-b)^2)=|y0-c|

OpenStudy (alexandervonhumboldt2):

Square both sides.

OpenStudy (sleepyhead314):

I am very confused by all this terminology @AlexandervonHumboldt2 you are typing very very fast

jaynator495 (jaynator495):

iambatman you still havent answered my question ._.

OpenStudy (sleepyjess):

*screams head off*

OpenStudy (sleepyjess):

*rips hair out*

OpenStudy (alexandervonhumboldt2):

Expand the expression in y0 on both sides and simplify.

OpenStudy (sleepyjess):

I'm so confused with all of this :/

jaynator495 (jaynator495):

funny how im stalking batman... very ironic xD

OpenStudy (alexandervonhumboldt2):

ok here it summaries up

OpenStudy (nikato):

 sketch of what your graph would look like